OpenStudy (anonymous):

Solve the equation using factoring techniques. x ( x + 8 ) = 713

hero (hero):

Can be solved using factoring techniques but it would take more effort than what I am willing to do.

OpenStudy (anonymous):

I don't understand how to do the factoring techniques but I know that is what I have to use. It tells me in the question.

hero (hero):

If you're not familiar with factoring techniques, then you shouldn't be doing this question. It is way more advanced for you.

hero (hero):

Are you sure you posted it correctly?

OpenStudy (anonymous):

It is a college algebra class and I am trying to understand these problems. This is one of the problems on my exercise.

OpenStudy (anonymous):

and yes. I posted it correctly.

Directrix (directrix):

x ( x + 8 ) = 713 x^2 + 8x = 713 x^2 + 8x - 713 = 0 (x - 23) (x + 31) = 0 Use the Zero Product Property x - 23 = 0 or x + 31 = 0 x = 23 or x = -31

hero (hero):

Hmmm, I assumed that 713 was not factorable